OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box.

OpenStudy (anonymous):

help plz!

OpenStudy (inowalst):

Okay. I'll help.

OpenStudy (anonymous):

Great!

OpenStudy (inowalst):

\[\frac{ 2 }{ ? }=\frac{ 25 }{ 60 }\]Multiply 60 times 2. Then divide that by 25.

OpenStudy (anonymous):

4.8

OpenStudy (inowalst):

That's your answer.\[\frac{ 2 }{ 4.8 }=\frac{ 25 }{ 60 }\]

OpenStudy (anonymous):

Thanks!

OpenStudy (inowalst):

You're welcome.

OpenStudy (anonymous):

so the answer is 4.8

OpenStudy (inowalst):

Yes.

OpenStudy (anonymous):

now can you help me on part b?

OpenStudy (inowalst):

Sure.